T1UNMAC(1) T1UNMAC(1) NAME t1unmac−translateaMacPostScriptfontintoPFAorPFBformat SYNOPSIS t1unmac[−a|−b][−r][input[output]] DESCRIPTION t1unmacextractsPOSTresourcesfromaMacintoshPostScriptfontfileandcreatesaPFA(hexadecimal)orPFB(binary)fontfile.ThefileinputshouldbeinMacBinaryIorII,AppleSingle, AppleDouble,orBinHexformat,oritcanbearawresourcefork.Ifthefileisarawresourcefork, youneedtogivethe‘−−raw’option;otherwiset1unmacshouldautomaticallyfigureoutwhat kindoffileyouhave.Ifthefileoutputisnotspecifiedoutputgoestothestandardoutput. OPTIONS −−pfa,−a OutputinPFA(ASCII)format. −−pfb,−b OutputinPFB(binary)format.Thisisthedefault. −−raw,−r Indicatesthattheinputisarawresourcefork. −−macbinary IndicatesthattheinputisinMacBinaryIorIIformat. −−applesingle IndicatesthattheinputisinAppleSingleformat. −−appledouble IndicatesthattheinputisinAppleDoubleformat. −−binhex IndicatesthattheinputisinBinHex4.0format. −−block−length=num,−lnum PFBonly:Setthemaximumoutputblocklengthtonum.Thedefaultlengthisaslargeas memoryallows. −−line−length=num,−lnum PFAonly:Setthemaximumlengthofencryptedlinesintheoutputtonum.(Thesearethe linesconsistingwhollyofhexadecimaldigits.)Thedefaultis64. SEEALSO t1mac(1),t1ascii(1),t1binary(1),t1asm(1),t1disasm(1) AUTHORS LeeHetherington(ilh@lcs.mit.edu) EddieKohler(eddietwo@lcs.mit.edu) PortedtoMicrosoftC/C++CompilerandMS-DOSoperatingsystembyKai-UweHerbing(herb-ing@netmbx.netmbx.de). 5Mar1998 1